OpenStudy (anonymous):

If only 10 percent of the 40 math homework problems are multiple choice, how many problems are multiple choice? A. 2 B. 4 C. 10 D. 40

OpenStudy (anonymous):

Did you do it? What's your answer?

OpenStudy (anonymous):

\[\frac{ 10 }{ 100 } \times 40\] Can you do it now?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

can you do that to this If 15 percent of the sixth graders have blonde hair, and there are 120 sixth graders, how many of them have blonde hair? A. 8 B. 15 C. 18 D. 22 to

OpenStudy (anonymous):

can you do that to this one to ? @Jas9420

OpenStudy (anonymous):

\[\frac{ 15 }{ 100} \times 120 \] Can you solve it now?

OpenStudy (anonymous):

18 and can you do that to all my problems i just have a couple =)

OpenStudy (anonymous):

@Jas9420

OpenStudy (anonymous):

The attendance last year at the Rock Brothers Music Festival was 567 people. This year the attendance was 670 people. What is the percent of increase for this year? Round to the nearest whole percent. A. 18% B. 20% C. 15% D. 19% @Jas9420

OpenStudy (anonymous):

A. 670-567=103 103/567=?/100 cross multiply 100*103=10300/567=18 ^